# On the universality of the halo mass function beyond $\Lambda$CDM cosmology

### v1

@@ 6 Conclusions
In order to test the universality $\nu f(\nu)$ we built a linear interpolation model of the results from our $z=0$ fiducial run. On comparison of the $\nu f(\nu)$ measured from the variational runs with this model, we were able to assess the validity of the putative universal behaviour with respect to the 16 variations. We also selected a number of snapshots with a coverage of redshift $z\in[0,6.75]$ to inspect evolution with redshift. To make quantitative statements, we performed a reduced-$\chi^{2}$ test, for which we applied the subcubing approach introduced in Smith & Marian (2011) to compute the covariance matrices of the HMFs and so estimate statistical errors. We summarise our findings as follows:
@@
In recent years there has been a rapid increase in the methods of identifying and classifying dark matter haloes in numerical simulations and in observational data. A number of theoretical studies have followed the path of studying haloes identified through the SO algorithm. However, it has been widely confirmed that the HMFs for these objects are far from a universality and thus require a large amount of simulations to accurately calibrate their abundance as a function of cosmological model and over time. By contrast, haloes identified through the FoF algorithm exhibit approximate universality with respect to a wide set of cosmological parameters. The advantage of this is that only several mass functions need to be accurately parameterised so that they can be deployed for likelihood modelling in cluster abundance studies.
In this work, we have completed a first step to producing an improved model for the HMF. We have quantitatively shown that FoF haloes do have the potential for building such models with the minimum required sets of numerical simulations. Our next step will be searching for an effective form of the universal $\nu f(\nu)$ that can provide percent-level constraints. The ultimate goal, of course, is to produce a theory forward modelling pipeline for modelling the likelihood function for observing cluster counts binned in terms of observable mass proxies for upcoming galaxy surveys such as Euclid. To achieve this, one needs to understand the conditional probability density of obtaining the observable given the underlying halo.
@@ Acknowledgements
RES acknowledges support from the Science and Technology Facilities Council (grant ST/X001040/1). YL thanks Baojiu Li for useful discussion. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure. We acknowledge that the results of this research have been achieved using the PRACE Tier-0 Research Infrastructure resource SuperMuc based in Garching Germany at the Leibniz Supercomputing Centre (LRZ) under project number 2012071360.
@@
Appendix A Effect of binning on multiplicity function
We quantified the systematic bias on the binned mass function regarding bin width. Firstly, we assume some forms of the fitting functions as true mass functions, which are obtained by a HMF calculator hmf (Murray et al., 2013). Then we adopt our fiducial cosmology to estimate the binned mass functions from the fitting functions by computing the numerical integral of each bin over its bin width. The relative differences between the binned $\nu f(\nu)$ and the fitting $\nu f(\nu)$ are shown in Fig. 9, which indicates that binning would cause an overestimate of $\nu f(\nu)$. This can be seen with all 3 selected models of fitting functions, i.e., Watson(FoF) (squares), Tinker08 (triangles), and Bhattacharya (circles), although the extent of the overestimate could be a little different regarding each model. The left, middle and right panels are for a choice of 5, 8, and 11 bins per decade, respectively, in logarithmic scale of halo masses. By comparing them, it is understood that the bias can be theoretically reduced by increasing the number of bins applied. However, in practice, small bin width could cause a problem of empty bins as well as very wiggled mass function when counting haloes in numerical simulations. Thus an optimal choice of bin width should be checked case by case. Additionally, it can be seen that the binned mass functions tend to be much more biased towards high redshift with the same choice of bin width. For example, an overestimate is below $2\%$ at $z=0$ with 8 bins per decade used, whereas it could reach a maximum of $6\%$ at $z=6.75$. Therefore, if one tries to measure an actual mass function with binning method, the correction of the bias respecting bin width should be seriously taken into account.
@@ Appendix A Effect of binning on multiplicity function
Appendix B Results of correlation matrices for individual runs of fiducial cosmology
@@
Appendix C Results of multiplicity function for higher redshifts
Fig. 11 shows the results of $\nu f(\nu)$ at $0\leq z\leq 6.75$. The assumed universal loci are wiggled at low redshift and tend to be smooth towards high redshift. It can be seen that the $\nu f(\nu)$ measured from all models are very close to the universal locus at each of the redshifts, especially going to high redshift.

### v2

@@ 6 Conclusions
In order to test the universality $\nu f(\nu)$ we built a linear interpolation model of the results from our fiducial run. On comparison of the $\nu f(\nu)$ measured from the variational runs with this model, we were able to assess the validity of the putative universal behaviour with respect to the 16 variations. We also selected a number of snapshots with a coverage of redshift $z\in[0,6.75]$ to inspect evolution with redshift. To make quantitative statements, we performed a reduced-$\chi^{2}$ test, for which we applied the subcubing approach introduced in Smith & Marian (2011) to compute the covariance matrices of the HMFs and so estimate statistical errors. We summarise our findings as follows:
@@
In recent years there has been a rapid increase in the methods of identifying and classifying dark matter haloes in numerical simulations and in observational data. A number of theoretical studies have followed the path of studying haloes identified through the SO algorithm. However, it has been widely confirmed that the HMFs for these objects are far from a universality except using the virial overdensity and thus require a large amount of simulations to accurately calibrate their abundance as a function of cosmological model and over time. By contrast, haloes identified through the FoF algorithm exhibit approximate universality with respect to a wide set of cosmological parameters. The advantage of this is that only several mass functions need to be accurately parameterised so that they can be deployed for likelihood modelling in cluster abundance studies.
In this work, we have completed a first step to producing an improved model for the HMF. We have quantitatively shown that FoF haloes do have the potential for building such models with the minimum required sets of numerical simulations. Our next step will be searching for an effective form of the universal $\nu f(\nu)$ that can provide percent-level constraints. The ultimate goal, of course, is to produce a theory forward modelling pipeline for use in the likelihood function for observing cluster counts binned in terms of observable mass proxies for upcoming galaxy surveys such as Euclid. To achieve this, one needs to understand the conditional probability density of obtaining the observable given the underlying halo.
@@ Acknowledgements
We would like to thank Adrian Jenkins and Ravi Sheth for their careful reading of the manuscript and useful suggestions, which have improved the paper. RES acknowledges support from the Science and Technology Facilities Council (grant ST/X001040/1). YL thanks Baojiu Li for useful discussion. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure. We acknowledge that the results of this research have been achieved using the PRACE Tier-0 Research Infrastructure resource SuperMuc based in Garching Germany at the Leibniz Supercomputing Centre (LRZ) under project number 2012071360.
@@
Appendix A Effect of the Fourier lattice on the mass variance
Figure 10: Top panel: Variance of the matter fluctuations as a function of halo mass at $z=0$. The solid orange line shows an idealised reference base from computing $\sigma(M)$ as given in Eq. (15), using a continuum of Fourier modes with a large-scale cutoff set at the Horizon scale $k_{\rm min}=10^{-4}\,h\,{\rm Mpc}^{-1}$ and a small-scale cutoff of $k_{\rm max}=10^{2}\,h\,{\rm Mpc}^{-1}$. The dashed lines shows the variance computed using only the linear power spectrum, evolved to the present day that is sampled on the Fourier lattice when generating the Gaussian random field for the initial conditions, with different box size $L$ and grid size $N$. The vertical dotted lines correspond to $M_{\rm Ny}$, the mass-scale associated with haloes whose linear theory radius is given by the Nyquist scale of the initial Fourier mesh. Bottom panel: fractional difference of the variance measures with respect to the continuum case. The line-styles are as above.
In our simulation setup, we generate a Gaussian random field of fluctuations defined by the linear matter power spectrum, which is realised on a finite Fourier lattice using a PM grid of size $N_{\rm g}=2048^{3}$. For our simulation box, this corresponds to a fundamental mode $k_{\rm f}=2\pi/L=0.0126\,\,h\,{\rm Mpc}^{-1}$ and a Nyquist frequency $k_{\rm Ny}=\pi N_{\rm g}/L=12.9\,\,h\,{\rm Mpc}^{-1}$. To assess the impact of missing variance from $k$-space discreteness, we follow and extend the method of Smith et al. (2003) to estimate the $\sigma(M)$ relation on the lattice as follows:
$\sigma^{2}_{\rm lat}(M)$ $=\sum_{lmn}\frac{(\Delta k)^{3}}{(2\pi)^{3}}P(k_{lmn})W^{2}(k_{lmn}R)$
$=\frac{2\pi^{2}}{V}\sum_{l=-N_{\rm g}/2}^{N_{\rm g}/2}\sum_{m=-N_{\rm g}/2}^{N_{\rm g}/2}\sum_{n=-N_{\rm g}/2}^{N_{\rm g}/2}\frac{\Delta^{2}(k_{lmn})}{k_{lmn}^{3}}W^{2}(k_{lmn}R)$
$=\frac{16\pi^{2}}{V}\sum_{l=0}^{N_{\rm g}/2}\sum_{m=0}^{N_{\rm g}/2}\sum_{n=0}^{N_{\rm g}/2}\frac{\Delta^{2}(k_{lmn})}{k_{lmn}^{3}}W^{2}(k_{lmn}R)\ ,$ (26)
where $k_{lmn}\equiv 2\pi/L(l^{2}+m^{2}+n^{2})^{1/2}$ and $\Delta^{2}(k)\equiv 4\pi k^{3}P(k)/(2\pi)^{3}$. Note that in the second line the upper and lower limit of the sums is given by the Nyquist frequency normalised by the fundamental mode. The final line uses the fact that the power spectrum and filter scale depend only on the magnitude of the $k$-vector and so the inherent symmetry means that we only need to compute the sum over one octant. In practice, we individually compute the variance of the interior $k$-points, the planes and axes, and then multiply by their corresponding weights. The others are identical, either due to symmetry, rotation or parity invariance. Also note that $\lim_{k\rightarrow 0}\Delta^{2}(k)/k^{3}\rightarrow 0$, for CDM-like power spectra. This enables the lattice computation to be sped up by an order of magnitude. Finally, we note that this would not be true if we instead used the actual realisation of the power spectrum for the initial conditions, since in this case we would only have the symmetry $\delta^{*}({\bf k})=\delta(-{\bf k})$, for a real field, which means that only the upper half of $k$-space is independent.
Figure 10 shows that the results computed using the Fourier lattice agree to high precision for $M\gtrsim M_{\rm Ny}$, where we define the Nyquist mass scale as:
$M_{\rm Ny}$ $=\frac{4}{3}\pi\bar{\rho}\left(\frac{2\pi}{k_{\rm Ny}}\right)^{3}=\frac{4}{3}\pi\bar{\rho}\left(\frac{2\pi L}{\pi N_{\rm g}}\right)^{3}=\frac{32}{3}\pi m_{\rm p}\left(\frac{N_{\rm part}}{N_{\rm g}}\right)^{3}\ ,$ (27)
which corresponds to the range where the haloes are resolved by more than $\sim 33$ particles. Given that the haloes we considered in this work are sampled by at least 150 particles, the impact of the missing variance owing to the Fourier lattice can be considered to be negligible.
Appendix B Effect of binning on multiplicity function
@@
We quantified the systematic bias on the binned mass function regarding bin width. Firstly, we assume some forms of the fitting functions as true mass functions, which are obtained by a HMF calculator hmf (Murray et al., 2013). Then we adopt our fiducial cosmology to estimate the binned mass functions from the fitting functions by computing the numerical integral of each bin over its bin width. The relative differences between the binned $\nu f(\nu)$ and the fitting $\nu f(\nu)$ are shown in Fig. 11, which indicates that binning would cause an overestimate of $\nu f(\nu)$. This can be seen with all 3 selected models of fitting functions, i.e., Watson(FoF) (squares), Tinker08 (triangles), and Bhattacharya (circles), although the extent of the overestimate could be a little different regarding each model. The left, middle and right panels are for a choice of 5, 8, and 11 bins per decade, respectively, in logarithmic scale of halo masses. By comparing them, it is understood that the bias can be theoretically reduced by increasing the number of bins applied. However, in practice, small bin width could cause a problem of empty bins as well as very fluctuating mass function when counting haloes in numerical simulations. Thus an optimal choice of bin width should be checked case by case. Additionally, it can be seen that the binned mass functions tend to be much more biased towards high redshift with the same choice of bin width. For example, an overestimate is below $2\%$ at $z=0$ with 8 bins per decade used, whereas it could reach a maximum of $6\%$ at $z=6.75$. Therefore, if one tries to measure an actual mass function with binning method, the correction of the bias respecting bin width should be seriously taken into account.
Appendix C Results of multiplicity function for individual runs of fiducial cosmology
Fig. 12 shows a comparison of the measured $\nu f(\nu)$ for 10 individual runs and the mean of them in the fiducial cosmology at $z=0$. Each individual realisation exhibits distinct features, particularly for $\nu\gtrsim 2$, where deviations from the mean locus gradually exceeds $2\%$. However, when averaging over all 10 runs, the resulting $\nu f(\nu)$ locus appears significantly smoother.
Figure 12: Top panel: Comparison of the measured $\nu f(\nu)$ as a function of $\nu$ for 10 individual runs (coloured lines) and the mean of them (black line) in the fiducial cosmology at $z=0$. Bottom panel: Fractional difference of the data of each run relative to the mean. The grey area marks the relative difference of ±$\pm 2\%$. Lines are as before.
Appendix D Results of correlation matrices for individual runs of fiducial cosmology
@@
Appendix E Results of multiplicity function for higher redshifts
Fig. 14 shows the results of $\nu f(\nu)$ at $0\leq z\leq 6.75$. The assumed universal loci are fluctuating at low redshift and tend to be smooth towards high redshift. It can be seen that the $\nu f(\nu)$ measured from all models are very close to the universal locus at each of the redshifts, especially going to high redshift.